Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Proba

Expert replies
Source: — Problem Solving |

by [email protected] » Thu Mar 10, 2016 4:28 pm
Hi didieravoaka,

When dealing with probability questions, there are only two results that can be calculated: what you WANT to have happen OR what you DON'T WANT to have happen. Those two outcomes create the following equation:

(Want) + (Don't Want) = 1

Sometimes it's actually easier to calculate what you WANT to have happen by calculating what you DON'T WANT to have happen (and then subtract that fraction from the number 1).

Here, we're asked for the probability of flipping AT LEAST one head and AT LEAST one tail on 6 coin flips. Since each coin flip has 2 possible outcomes, there are 2^6 = 64 possible outcomes (although there would be lots of 'duplicate results'). We don't want to have to determine every possible outcome that gives us at least 1 head and at least one tail though, so let's calculate the probability of that NOT happening.

There are two results that would NOT fit what we're looking for:

ALL HEADS
ALL TAILS

The probability of each is the same: 1/64

1/64 + 1/64 = 2/64 = 1/32

1/32 = about 3%

(Want) + (.03) = 1

Want = about 97%

Final Answer: E

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by didieravoaka » Mon Mar 14, 2016 12:27 pm
Thanks Rich!

Can you detail the probability that NOT happening? Is that what gives 1/64?

Is that this way of thinking you apply... P(At least 1)=1-P(None)?

Thanks.
Join the discussion

by [email protected] » Mon Mar 14, 2016 9:41 pm
Hi didieravoaka,

When a probability question asks for something happening 'at least' a certain number of times, it's often easier to calculate what you DON'T WANT to have happen (in this case, the situations would NOT include at least one heads and one tails).

Here, the two options that would NOT fit what we're looking for are the "ALL HEADS" and "ALL TAILS" options. In all other situations, we'd have at least one heads and at least one tails. Since the "ALL HEADS" and "ALL TAILS" options are fairly easy to calculate, THAT method is arguably the fastest way to get to the correct answer.

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Matt@VeritasPrep » Thu Mar 17, 2016 10:12 pm
didieravoaka wrote:Thanks Rich!

Can you detail the probability that NOT happening? Is that what gives 1/64?

Is that this way of thinking you apply... P(At least 1)=1-P(None)?

Thanks.
That equation is right, and it's a GREAT one to use in any problem that asks you for the probability of "at least one" thing happening: it tends to save a LOT of time.
Join the discussion